# Building Information Modelling MRR modular reconfigurable robot

### v1

@@ Abstract
In situ robotic automation in construction is challenging due to constantly changing environments, a shortage of robotic experts, and a lack of standardized frameworks bridging robotics and construction practices. This work proposes a holistic framework for construction task specification, optimization of robot morphology, and mission execution using a mobile modular reconfigurable robot. Users can specify and monitor the desired robot behavior through a graphical interface. Our framework identifies an optimized robot morphology and enables automatic real-world execution by integrating Building Information Modelling (BIM). By leveraging modular robot components, we ensure seamless and fast adaption to the specific demands of the construction task. Experimental validation demonstrates that our approach robustly enables the autonomous execution of robotic drilling.
@@
This work addresses the challenge of automating robotic drilling on construction sites and introduces a general framework that also supports other construction tasks. Construction sites are unpredictable and constantly changing, making automation difficult. To tackle this, we propose a modular robotic system that can adapt to specific tasks. We use Building Information Modelling (BIM) data to automatically determine the best robot configuration by optimizing factors such as setup time and robustness to positioning errors. Through an intuitive interface, workers can specify tasks, after which they assemble the robot following the provided instructions.
This approach enables the same hardware to be deployed across various scenarios while accounting for site-specific challenges. The core innovation lies in combining modular robotics with automated optimization, task planning, and motion planning to minimize the need for user intervention while delivering robots specifically tailored to each task. To achieve this, the work extends existing methods for modular robot design, enabling consideration of multiple, potentially conflicting objectives in the optimization process. Experimental results demonstrate the ability of the system to perform autonomous drilling with high precision. While the system requires initial calibration, future improvements in vision and force control could further enhance its accuracy and robustness. Beyond drilling, this framework has the potential to support a wide range of applications, such as welding, painting, and assembly, offering a practical and adaptable solution for diverse construction tasks.
@@
Robotic Driling, Modular Robots, BIM
@@ I Introduction
Figure 1: Section A shows the available robot modules for the manipulator, composed of six joints (green), three static links (silver), and a custom end-effector. Section B shows the assembled robotic arm on a mobile base and the experimental setup.
In recent decades, robotic automation in manufacturing significantly increased productivity [1]. To this day, a comparative technology leap in the construction sector has not occurred [2, 3, 4, 5]. Additionally, harsh, non-ergonomic working conditions result in labor shortages in the construction industry [6]. Robotic automation in construction, however, requires versatility, adaptability, and robustness as construction sites are complex, constantly changing, potentially hazardous, and hard to navigate [7].
The main contribution of this paper is a framework for in situ robotic automation that features (i) an autonomous decomposition of high-level instructions (e.g., “drill holes” or “paint wall”) into executable skills, (ii) identification of an optimized module composition for the robot arm using BIM data of the environment, (iii) online motion planning, and (iv) adaptions to positioning errors of the mobile base. The only actions required by the human worker are specification of the task and assembly of robot modules once the optimization of their composition terminates. While our framework is not limited to a particular task, we focus on autonomous drilling as shown in  Figure 1. A supplementary video of the real-world experiment and more can be found on the project website modular-construction-robot.cps.cit.tum.de.
@@ II Related Work
This section discusses previous applications of mobile manipulators in construction and, specifically, drilling use cases. Further, we provide an overview of current methods for optimizing modular robot compositions.
@@ II-A Robotic automation in building construction
Some construction tasks, such as cutting bricks or insulation panels, can be automated by special machinery or onsite manufacturing processes [8]. However, highly accurate and flexible interactions with the environment necessitate mobile robotic systems that can self-localize [9]. Navigating the highly unstructured environments of construction sites challenges the sensing and perception capabilities of robots. In [10], 3D scans taken with onboard sensors were used to enable navigation and localization of a bricklaying robot in situ. The perception can be further enriched with data from as-built BIM [11].
Leveraging BIM data in construction has been shown to increase worker productivity [12] and accelerate the deployment of construction robotics [13]. However, as reported by [14], there is still a lack of research on integrating BIM with onsite robotic systems. Previous work proposed the integration of BIM data into a mission parametrization toolchain for non-experts and validated its effectiveness through a paint-spraying experiment [15].
Particularly challenging use cases in construction robotics are precise power tool works on walls or ceilings, such as grinding or drilling holes. The challenges stem from uncertainties in robot perception and disturbances induced by operating the power tools. Recent works present an aerial system for drilling holes limited to free space areas and vertical walls or facades [16] or the development of a mobile manipulator with predefined reach and dexterity [17].
Robots with a fixed kinematic, as presented by industrial manufacturers111for instance, “Hilti Jaibot”, “Baubot”, or “Schindler R.I.S.E” or in previous research [18, 19, 20], may not provide the necessary flexibility required in construction [21]. Reconfigurable modular robots – comprising modular joints, links, and sensor modules – offer better transportability to and on the construction site and can also be adapted to specific requirements in situ. This aligns with previous research on digital in situ fabrication [22], which highlights that versatile, modular, and reusable solutions and real-world experiments are required to enable innovation in construction automation.
@@ II-B Optimizing modular robot composition
The optimization of robot morphologies, and more specifically, finding an optimal module composition for reconfigurable robots, has attracted attention for decades [23, 24, 25, 26, 27]. Some approaches introduce heuristics and surrogate performance criteria to enable a brute-force search over reduced search spaces [28, 29, 30]. Most approaches, however, use discrete black-box optimization algorithms [31, 32, 33]. Since the early work of Karl Sims [34], genetic algorithms (GAs) have been a common choice for the optimization of modular robot morphologies. The work in [35] deploys a genetic algorithm (GA) to co-optimize the morphology and base placement of a modular robot. A lexicographic genetic approach is proposed in [36] to increase the efficiency and interpretability of the evolutionary optimization process. While our work builds upon [36], we extend it to multi-objective problems to explicitly model the tradeoff between performance and robustness. Further, in contrast to previous work, we perform an experimental validation of the optimized designs.
II-C Robustness and multi-objective optimization
It is well-known that a purely simulation-based optimization of real-world systems can lead to suboptimal performance during deployment – a phenomenon frequently termed the reality gap [37]. This is particularly problematic in unstructured or partially structured environments, like construction sites, where conditions are highly dynamic and unpredictable, and simulation often fails to capture the full complexity of real-world variability. Especially in learning-based approaches, domain randomization has been successfully applied to tackle this issue [38, 39, 40]. However, past works indicate a tradeoff between the specificity and generality of a solution [41, 42]. In this work, we explicitly model this tradeoff by formulating a multi-objective optimization problem with performance and robustness against positioning errors as competing objectives to find Pareto-optimal solutions. To this end, we incorporate the well-known nondominated sorting-based genetic algorithm (NSGA-II[43] into the lexicographic optimization procedure previously introduced in [36].
@@ III Framework
Figure 2: The structure of the proposed framework from optimization of a robot model to mission execution. During mission preparation, inputs from the user interface are used to perform robot synthesis. The mission execution and robot control block run during deployment.
Our framework is shown in Figure 2. It can be separated into four modules: The user interface and robot synthesis modules are required for mission preparation, whereas mission execution and robot control run during deployment. A user generates and parameterizes a mission through the BIM-based user interface presented in [15]. Based on the user input, the robot synthesis module searches for the optimal configuration of the robot modules in simulation to execute the provided mission and proposes it to the user. Next, the user assembles the modular robot and activates its control setup. Once the robot is ready, the initially configured mission is sent to the robot. Finally, the robot control module adapts an initially planned trajectory based on visual calibration and executes it on the real robot.
@@ III-A Robot Synthesis
Reconfigurable modular robot compositions are simulated and optimized with Timor [44]. We rely on the modular robot modeling language introduced in CoBRA [45] and a custom parser module (CoBRA BT Parser) to integrate the user interface and the BIM model with Timor. After the optimization algorithm described in Section IV identifies an optimized module composition and a motion trajectory, the worker can proceed with the physical robot assembly. The self-identification of the robot [46] or automatic model generation [44] can be used to obtain a kinematic and dynamic model of the robot once its modules are provided. The motion trajectory is stored and later passed to the robot control module.
@@ III-B Robot Control
The robot control follows a modular setup. While our framework is agnostic to the specific software being used for control, we integrated the following components for deployment:
  • The mobile base navigation, including local and global path planning, localization, and omnidirectional control, is based on the navigation2 framework [47, 48].
  • The manipulator control, including collision-free trajectory planning and execution, is based on Timor and joint impedance control using the manipulator control library [^2] https://github.com/FraunhoferItalia/fhi_manipulator_control_stack. .
  • The calibration module based on OpenCV [^3] https://github.com/opencv/opencv includes a pipeline that detects ArUco landmark poses [49] with a camera to calibrate the robot position with respect to the environment. As the mobile base cannot move with arbitrary precision, even after calibration, the manipulator might not exactly be in the desired position – in this paper, we refer to this remaining error as the base positioning error.
@@ IV Modular Robot Composition Optimization
The following section introduces our method for optimizing modular robot compositions given a task description and assumptions about the positioning accuracy of the mobile base. Throughout this section, we follow the notation shown in Table II.
@@ IV-A Preliminaries
As a running example, consider a ranking of robots $\{1,2,3,4\}$, based on how fast they can perform a specific task ($f_{fast}$), how cheap the corresponding hardware is ($f_{cheap}$), and how robust they are to external disturbances ($f_{robust}$). Table I shows the scores ${\mathbf{r}}_{i}$ of four exemplary robots. The higher the score, the better the robot. There are multiple objectives, so no clear “best” robot exists. In the following sections, we will introduce a ranking-based selection method that considers multiple objectives.
@@
Let $(f_{fast},f_{cheap},f_{robust})$ be the sequence of lexicographic fitness values for the exemplary robots. As the fourth robot is the fastest one and robot three is cheaper than robot two, while they are equally fast, we obtain the lexicographic order ${\mathbf{r}}_{4}>{\mathbf{r}}_{3}>{\mathbf{r}}_{2}>{\mathbf{r}}_{1}$. However, if we chose our fitness sequence to be $(f_{robust},f_{fast},f_{cheap})$, the order would be ${\mathbf{r}}_{1}>{\mathbf{r}}_{4}>{\mathbf{r}}_{3}>{\mathbf{r}}_{2}$.
@@
Let $(f_{fast},f_{cheap},f_{robust})$ be a multi-objective fitness sequence for the exemplary robots. We can see that ${\mathbf{r}}_{3}{\succ}{\mathbf{r}}_{2}$ as both robots are equally fast and robust, but robot three is cheaper. However, robots one, three, and four are all Pareto-optimal as none of them is dominated by any other robot.
Genetic algorithm. Genetic algorithms change the population of different solution candidates (individuals) to maximize an objective, called the fitness function [53]. Every individual is represented by a genome that encodes its unique attributes. For a fixed number of iterations, a new population is created by selecting a limited number of individuals from the previous population (selection), creating offsprings by combining attributes of existing individuals (crossover), and mutating some of the genomes (mutation). A cycle of selection, crossover, and mutation is called a generation.
NSGA-II. We write $\mathrm{ndr}({\mathbf{a}},{\mathbb{A}})$ and $\mathrm{cd}\left({\mathbf{a}},{\mathbb{A}}\right)$ whenever we compute the nondomination rank or the crowding distance of an individual ${\mathbf{a}}$ with respect to a population ${\mathbb{A}}$ according to the NSGA-II algorithm as proposed in [43].  -A provides an informal introduction to these metrics.
@@ IV-B Lexicographic nondominant sorting optimization
$\left[n\right]$ The set $\{1,2,\dots,n\}$ $({\mathbb{A}},\leq)$ Partially ordered sets
$\mathrm{Tr}({\bm{M}})$ The trace of a matrix $|{\mathbb{A}}|$ The cardinality of a set
@@
$\mathrm{cd}\left({\mathbf{a}},{\mathbb{A}}\right)$ The crowding distance of a multi-objective value ${\mathbf{a}}$ with respect to its nondomination front in set ${\mathbb{A}}$
$\mathrm{ndr}({\mathbf{a}},{\mathbb{A}})$ The nondomination rank of a multi-objective value ${\mathbf{a}}$ with respect to the set ${\mathbb{A}}$
@@
We perform a mixed Pareto-lexicographic optimization based on the approach proposed in [36]. This work introduces a genetic algorithm with a selection that is based on a purely lexicographic fitness function. Notably, the approach initially targets a single-objective optimization. The introduction of a lexicographic fitness function is motivated by computational efficiency: By introducing multiple surrogate objectives, the selection process of the genetic optimization can be performed without evaluating the computationally expensive primary objective for a majority of solution candidates.
Our work combines this approach with the NSGA-II algorithm [43] to achieve computationally efficient multi-objective optimization. We compute two sequences ${\mathbf{c}}_{i}$ and ${\mathbf{x}}_{i}$ for every individual $i$ in a population. The sequence ${\mathbf{c}}_{i}$ contains $m$ values that measure compliance with $m$ constraints that can be easily evaluated, such as the sufficient size of a robot or the absence of self-collisions. We define these constraints so that the $j$th value of ${\mathbf{c}}_{i}$, $c_{i,j}$ is 00, if solution $i$ satisfies constraint $j$, and negative, otherwise. In addition to previous work, we introduce the sequence
@@
where $a_{i},b_{i},\dots$ can be arbitrary objectives, such as the monetary cost or robustness of an individual solution. We write ${\mathbb{X}}=\{{\mathbf{x}}_{1},\dots,{\mathbf{x}}_{n}\}$ and define the lexicographic fitness as the sequence
@@
By performing a lexicographic selection based on this fitness, we can efficiently rule out candidates that do not fulfill the constraints. In addition, the nondomniation rank fitness from  (4) allows us to perform multi-objective selection under consideration of ${\mathbf{x}}_{1},\dots,{\mathbf{x}}_{n}$. Instead of a single solution, our optimization results in a set of Pareto-optimal solutions from which an operator can pick.

### v2

# Building Information Modeling MRR modular reconfigurable robot

@@ Abstract
In situ robotic automation in construction is challenging due to constantly changing environments, a shortage of robotic experts, and a lack of standardized frameworks bridging robotics and construction practices. This work proposes a holistic framework for construction task specification, optimization of robot morphology, and mission execution using a mobile modular reconfigurable robot. Users can specify and monitor the desired robot behavior through a graphical interface. In contrast to existing, monolithic solutions, we automatically identify a new task-tailored robot for every task by integrating Building Information Modeling (BIM). Our framework leverages modular robot components that enable the fast adaption of robot hardware to the specific demands of the construction task. Other than previous works on modular robot optimization, we consider multiple competing objectives, which allow us to explicitly model the challenges of real-world transfer, such as calibration errors. We demonstrate our framework in simulation by optimizing robots for drilling and spray painting. Finally, experimental validation demonstrates that our approach robustly enables the autonomous execution of robotic drilling.
@@
Construction sites are unpredictable and constantly changing, making automation difficult. To tackle this, we propose a modular robotic system that can adapt to specific tasks. In this work, we address the challenge of automating robotic drilling and introduce a general framework that also supports other construction tasks. We use Building Information Modelling (BIM) data to automatically determine the best robot configuration by optimizing factors such as setup time and robustness to positioning errors. Through an intuitive interface, workers can specify tasks, after which they assemble the robot following the provided instructions.
The result of this work is a robotic system that can be deployed on diverse construction sites while requiring only readily available BIM data and minimal user inputs. Our approach enables the same hardware to be deployed across various scenarios while accounting for site-specific challenges. A core innovation lies in combining modular robotics with automated optimization, task planning, and motion planning to minimize the need for user intervention while delivering robots specifically tailored to each task. To achieve this, we extend methods for modular robot design to account for multiple conflicting objectives in the optimization process. Experimental results demonstrate the ability of the system to perform autonomous drilling with high precision. Beyond drilling, this framework has the potential to support a wide range of applications, and we show its applicability to spray painting in simulation. Overall, it offers a practical and adaptable solution for diverse construction tasks and sites.
@@
Modular Robot, BIM, Construction Automation, Morphology Optimization
@@ I Introduction
Figure 1: Section A shows the available robot modules for the manipulator, composed of six joints (green), three passive links (silver), and a custom end effector. Section B shows the assembled robotic arm on a mobile base and the experimental setup.
In recent decades, robotic automation in manufacturing significantly increased productivity [1]. To this day, a comparative technology leap in the construction sector has not occurred [2, 3, 4, 5]. Furthermore, harsh and non-ergonomic working conditions result in labor shortages in the construction industry [6]. Addressing these shortages through robotic automation requires versatility, adaptability, and robustness, as construction sites are complex, constantly changing, potentially hazardous, and challenging to navigate [7].
In contrast to existing robotic systems for construction, the modular robot deployed in this work can easily be adapted to new tasks or changes in the environment in minutes. The main contribution of this paper is a framework for in situ robotic automation that features (i) an autonomous decomposition of high-level instructions (e.g., “drill holes” or “paint wall”) into executable skills, (ii) identification of an optimized module composition for the robot arm using BIM data of the environment, (iii) online motion planning, and (iv) adaptations to positioning errors of the mobile base. The only actions required by the human worker are the specification of the task and the assembly of robot modules once the optimization of their composition is complete. Although our framework is not limited to a particular task, for real-world evaluation, we focus on autonomous drilling as shown in Figure 1. A supplementary video of the real-world experiment and more can be found on the project website modular-construction-robot.cps.cit.tum.de.
@@ II Related Work
This section discusses previous applications of mobile manipulators in construction and, specifically, drilling use cases. In addition, we provide an overview of current methods for optimizing modular robot compositions. We conclude with an overview of common strategies to tackle challenges in the transfer of simulation to real-world application.
@@ II-A Robotic automation in building construction
Some construction tasks, such as cutting bricks or insulation panels, can be automated by special machinery or in situ manufacturing processes [8]. However, highly accurate and flexible interactions with the environment require mobile robotic systems that can self-localize [9]. Navigating the highly unstructured environments of construction sites challenges the sensing and perception capabilities of robots. In [10], 3D scans taken with onboard sensors enabled navigation and localization of a bricklaying robot in situ. Perception can also be enriched with data from as-built BIM [11].
The use of BIM data in construction has been shown to increase worker productivity [12] and accelerate the deployment of construction robotics [13]. However, as reported by [14], there is still a lack of research on integrating BIM with robotic systems in situ. Previous work proposed the integration of BIM data into a mission parameterization toolchain for non-experts and validated its effectiveness through a spray-painting experiment [15].
Particularly challenging use cases in construction robotics are precise power tool works on walls or ceilings, such as grinding or drilling holes. The challenges stem from uncertainties in robot perception and disturbances induced by the operation of the power tools. Recent works present an aerial system for drilling holes limited to free-space areas and vertical walls or facades [16] or the development of a mobile manipulator with predefined reach and dexterity [17].
Robots with fixed kinematics as presented in previous research [18, 19, 20], may not provide the necessary flexibility required in construction [21]. This limitation is shared by most existing industrial solutions: Robots such as “Baubot” or “Hilti Jaibot” can perform semi-autonomous drilling, but are limited to mostly unobstructed environments and predetermined operating heights. Others, such as the “Schindler R.I.S.E”, are tailored to specific environments, such as elevator shafts, but cannot be deployed anywhere else. Reconfigurable modular robots – comprising modular joints, links, and sensor modules – can be adapted to specific requirements in situ while offering better transportability to and on the construction site. This aligns with previous research on digital in situ fabrication [22], highlighting that versatile, modular, and reusable solutions and real-world experiments are required to enable innovation in construction automation.
@@ II-B Optimizing modular robot composition
The optimization of robot morphologies, and more specifically, finding an optimal module composition for reconfigurable robots, has attracted attention for decades [23, 24, 25, 26, 27]. Some approaches introduce heuristics and surrogate performance criteria to enable a brute-force search over reduced search spaces [28, 29, 30]. Most approaches, however, use discrete black-box optimization algorithms [31, 32, 33]. Since the early work of Karl Sims [34], genetic algorithms (GAs) have been a common choice for the optimization of modular robot morphologies. The work in [35] deploys a genetic algorithm (GA) to co-optimize the morphology and base placement of a modular robot. In  [36], a genetic lexicographic algorithm is proposed to increase the efficiency and interpretability of the evolutionary optimization process. While our work builds upon [36], we extend it to multiobjective problems to explicitly model the trade-off between performance and robustness. None of the above-mentioned works considers the challenges of deploying these robots outside idealized lab settings; the simulation is treated as a perfect representation of the world. In contrast, we perform robot drilling using modular hardware while considering the challenges of simulation-to-reality transfer, such as inaccurate positioning and visual calibration of the robot arm.
II-C Robustness and multiobjective optimization
It is well known that a purely simulation-based optimization of real-world systems can lead to suboptimal performance during deployment – a phenomenon commonly referred to as the reality gap [37]. This is particularly problematic in unstructured or partially structured environments, such as construction sites, where conditions are highly dynamic and unpredictable, and simulation often fails to capture the full complexity of real-world variability. Especially in learning-based approaches, domain randomization has been successfully applied to address this issue [38, 39, 40]. However, past work indicates a trade-off between the performance and generality of a solution [41, 42]. In this work, we explicitly model this trade-off by formulating a multiobjective optimization problem with performance and robustness against positioning errors as competing objectives to find Pareto-optimal solutions. To this end, we build on the nondominated sorting-based genetic algorithm (NSGA-II[43].
@@ III Framework
Figure 2: The structure of the proposed framework from optimization of a robot model to mission execution. The bold boxes indicate the contributions of this work. During mission preparation, inputs from the user interface are used to perform robot synthesis. The mission execution and robot control block run during deployment.
Our framework is shown in Figure 2. It can be separated into four modules: The user interface and robot synthesis modules are required for mission preparation, whereas mission execution and robot control run during deployment. A user generates and parameterizes a mission through the BIM-based user interface presented in [15]. Based on the user input, the robot synthesis module searches for Pareto-optimal configurations of the robot modules in simulation to execute the provided mission and proposes them to the user. Next, the user assembles the modular robot and activates its control setup. Finally, the robot performs visual calibration and executes the planned trajectory.
@@ III-A Robot Synthesis
Reconfigurable modular robot compositions are simulated and optimized using the toolbox for industrial modular robotics (Timor) [44]. We rely on the modular robot modeling language introduced in CoBRA [45] and a custom parser module (CoBRA BT Parser) to integrate the user interface and the BIM model using Timor. After the optimization algorithm described in Section IV identifies an optimized module composition and a motion trajectory, the worker can proceed with the physical robot assembly. Self-identification of the robot [46] or automatic model generation [44] can obtain a kinematic and dynamic robot model once its modules are provided. The motion trajectory is stored and later passed to the robot control module.
@@ III-B Robot control
The robot control follows a modular setup. While our framework is agnostic to the specific software that is used for control, we have integrated the following components for deployment:
  • Mobile base navigation, including local and global route planning, localization, and omnidirectional control, is based on the navigation2 framework [47, 48].
  • The manipulator control, including collision-free trajectory planning and execution, is based on Timor and joint impedance control using the manipulator control library [^1] https://github.com/FraunhoferItalia/fhi_manipulator_control_stack.
  • The calibration module based on OpenCV [^2] https://github.com/opencv/opencv includes a pipeline that detects ArUco landmark poses [49] with a camera to calibrate the robot position with respect to the environment. As the mobile base cannot move with arbitrary precision, even after calibration, the manipulator may not be exactly in the desired position. In this paper, we refer to this remaining error as the base positioning error.
@@ IV Modular Robot Composition Optimization
The modular nature of the robot realizes flexibly adapting its morphology. However, it is a challenging task even for human experts to find a suitable or even optimized configuration for a specific task [44]. In this section, we introduce an automatic optimization method for robot morphologies. In contrast to previous work on robot design, we consider and jointly optimize multiple, potentially competing objectives $(a,b,\dots)$, such as cycle time, robustness, or the monetary cost of a solution. Formally, the resulting constrained multiobjective optimization problem is given by
$\max_{{\textnormal{m}}_{i}\in{\mathbb{M}}}$ $\Bigl{(}a({\mathbf{m}},{\mathbb{G}},{\mathbb{O}}),b({\mathbf{m}},{\mathbb{G}},{\mathbb{O}}),\dots\Bigr{)}$ (1)
s.t. $c_{j}({\mathbf{m}},{\mathbb{G}},{\mathbb{O}})=0\quad\text{for }j\in[m]\,,$
where $c_{j}$ are task constraints to be considered, and a sequence of modules fully defines the robot morphology ${\mathbf{m}}=({\textnormal{m}}_{1},\dots,{\textnormal{m}}_{i})$. A summary of the notation can be found in Table II.
Both constraints and objectives can be evaluated based on robot morphology ${\mathbf{m}}$, task-specific end-effector goals ${\mathbb{G}}$, and obstacles ${\mathbb{O}}$. Without a clear hierarchy of objectives, (1) generally does not have a single optimal solution. Instead, our objective is to find all Pareto-optimal solutions. To this end, we introduce a method based on lexicographic nondominant sorting optimization that combines the efficiency of lexicographic genetic optimizers with the versatility of the NSGA-II algorithm. While lexicographic optimization is based on hierarchical ordering, we only use it to increase the efficiency of constraint evaluation for (1). No order is imposed on the optimization objectives $(a,b,\dots)$.
In Section IV-A, we include fundamental preliminaries, followed by formal definitions of our optimization approach, task definition, objective functions, and constraints in Sections IV-B to IV-D.
@@ IV-A Preliminaries
As a running example, consider a ranking of robots $\{1,2,3,4\}$, based on how fast they can perform a specific task ($f_{fast}$), how cheap the corresponding hardware is ($f_{cheap}$), and how robust they are to external disturbances ($f_{robust}$). Table I shows the scores ${\mathbf{r}}_{i}$ of four exemplary robots. The higher the score, the better the robot. There are multiple objectives, so there is no clear “best” robot. In the following sections, we will introduce a ranking-based selection method that considers multiple objectives.
@@
Let $(f_{fast},f_{cheap},f_{robust})$ be the sequence of lexicographic fitness values. As the fourth robot is the fastest one and robot three is cheaper than robot two, while they are equally fast, we obtain the lexicographic order ${\mathbf{r}}_{4}>{\mathbf{r}}_{3}>{\mathbf{r}}_{2}>{\mathbf{r}}_{1}$. However, if we chose the lexicographic order to be $(f_{robust},f_{fast},f_{cheap})$, the ranking would be ${\mathbf{r}}_{1}>{\mathbf{r}}_{4}>{\mathbf{r}}_{3}>{\mathbf{r}}_{2}$.
@@
Let $(f_{fast},f_{cheap},f_{robust})$ be a multiobjective fitness sequence for the exemplary robots. We can see that ${\mathbf{r}}_{3}{\succ}{\mathbf{r}}_{2}$ since both robots are equally fast and robust, but robot three is cheaper. However, robots one, three, and four are all Pareto-optimal as none is dominated by any other robot.
Genetic algorithm. Genetic algorithms change the population of different solution candidates (individuals) to maximize an objective, called the fitness function [53]. Each individual is represented by a genome that encodes its unique attributes. For a fixed number of iterations, a new population is created by selecting a limited number of individuals from the previous population (selection), creating offsprings by combining attributes of existing individuals (crossover), and mutating some of the genomes (mutation). A cycle of selection, crossover, and mutation produces a new generation.
NSGA-II. We write $\mathrm{ndr}({\mathbf{a}},{\mathbb{A}})$ and $\mathrm{cd}\left({\mathbf{a}},{\mathbb{A}}\right)$ whenever we compute the nondomination rank or the crowding distance of an individual ${\mathbf{a}}$ with respect to a population ${\mathbb{A}}$ according to the NSGA-II algorithm as proposed in [43]. Appendix -A provides an informal introduction to these metrics.
@@ IV-B Lexicographic nondominant sorting optimization
$\left[n\right]$ The set $\{1,2,\dots,n\}$ $({\mathbb{A}},\leq)$ Partially ordered set
@@
$\mathrm{cd}\left({\mathbf{a}},{\mathbb{A}}\right)$ The crowding distance of a multiobjective value ${\mathbf{a}}$ with respect to its nondomination front in set ${\mathbb{A}}$
$\mathrm{ndr}({\mathbf{a}},{\mathbb{A}})$ The nondomination rank of a multiobjective value ${\mathbf{a}}$ with respect to the set ${\mathbb{A}}$
@@
We perform a mixed Pareto-lexicographic optimization based on the method introduced in [36], which presents a single-objective genetic algorithm with a purely lexicographic fitness function. The introduction of a lexicographic fitness function is motivated by computational efficiency: By introducing multiple surrogate objectives, the selection process of the genetic optimization can be performed without evaluating the computationally expensive primary objective for most solution candidates.
For every individual $i$ in a population, we compute two sequences ${\mathbf{c}}_{i}$ and ${\mathbf{x}}_{i}$ as follows: The sequence ${\mathbf{c}}_{i}$ contains $m$ values that measure compliance with $m$ constraints. Crucially, some of them are introduced because they can easily be evaluated, such as the sufficient size of a robot or the absence of self-collisions. We define these constraints so that the $j$th value of ${\mathbf{c}}_{i}$, $c_{i,j}$ is 00, if the solution $i$ satisfies the constraint $j$, and negative, otherwise. Further, we introduce the sequence
@@
where $a_{i},b_{i},\dots$ are our optimization objectives, such as the monetary cost or robustness of an individual solution. We write ${\mathbb{X}}=\{{\mathbf{x}}_{1},\dots,{\mathbf{x}}_{n}\}$ and define the lexicographic fitness as the sequence
@@
By performing a lexicographic selection based on this fitness, we can efficiently rule out candidates that do not fulfill the constraints. Further, the lexicographic selection process allows one to pick candidates that are closer to fulfilling the constraints if not enough feasible solutions exist. In addition, the nondomniation rank fitness from  (5) allows us to perform multiobjective selection considering ${\mathbf{x}}_{1},\dots,{\mathbf{x}}_{n}$. Instead of a single solution, our optimization results in a set of Pareto-optimal solutions from which an operator can pick.